The amount of pollutants A entering the atmosphere is directly proportional to the population P living in an area. If 25000 tons of pollutants enter the atmosphere in an area with a population of 17750, how many tons of pollutants enter the atmosphere in Manila with a population of 2450500? ( need a complete solution) thanks

Answer by robertb(5830)

A = kP is the working equation, with k = constant of proportionality. It is equal to k = 25000/17750 = 100/71. Then A = 100*2450500/71 = 3451408.45 tons, to 2 decimal places.
Manigong Bagong Taon! (If you can understand that...)